A physician is looking at a meta-analysis of the effects of short sleep duration on hyperglycemia during pregnancy. Data from eight studies were utilized. The findings show women that slept less than 6 hours had an odds ratio 1.70 of having gestational diabetes than those who sleep more than 6 hours.   This data can be interpreted that in women who slept less than 6 hours per night.

  1. Had a relative risk of <1 of having gestational
  2. Short sleep duration is a cause of gestational diabetes
  3. Did not have a statistically significant difference in their risk for f=gestational diabetes because the confidence interval was too narrow
  4. Had a 1.7 greater chance of having gestational diabetes compared to those who slept more than 6 hours
Answer #1

odds ratio 1.70 means

odds=p/1-p=probability than an individual has gestational diabetes who sleep less than 6 hours/probability than an individual has gestational diabetes who sleep more than 6 hours

odds=1.7/1

increased frequency of exposure among cases

ANSWER;

Had a 1.7 greater chance of having gestational diabetes compared to those who slept more than 6 hour